# ul Koszykowa 75, 00-662 Warsaw, Poland

### v4

@@ 3. Hypergeometric function
1) We have after inserting (2.8) in the last step
$_{2}F_{1}\left(a,b;c;\frac{\left(1-x\right)}{2}\right)=\sum_{n\geq 0}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}}(-1)^{n}\frac{\left(x-1\right)^{n}}{2^{n}n!}$
$=\sum_{n\geq 0}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}n!}(-1)^{n}\sum_{j=0}^{n}\tilde{e}_{n,j}(d,f)J_{j}(x|d,f)$
$=\sum_{j\geq 0}(-1)^{j}J_{j}\left(x|d,f\right)\sum_{n\geq j}(-1)^{n-j}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}n!}\tilde{e}_{n,j}(d,f)$
$\sum_{j\geq 0}(-1)^{j}J_{j}\left(x|d,f\right)\sum_{n\geq j}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}n!}\frac{n!\left(f+j\right)^{(n-j)}}{\left(n-j\right)!\left(d+f+j-1\right)^{\left(j\right)}\left(d+f+2j\right)^{\left(n-j\right)}}.$
By the well-known formula (see [4] (4.14)), we write
$_{2}F_{1}\left(a,b;c;\frac{\left(1-x\right)}{2}\right)=\sum_{j\geq 0}J_{j}\left(-x|f,d\right)$
$\times\sum_{n\geq j}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}}\frac{\left(f+j\right)^{(n-j)}}{\left(n-j\right)!\left(d+f+j-1\right)^{\left(j\right)}\left(d+f+2j\right)^{\left(n-j\right)}}.$
We now change x-x to $x$ and get
$_{2}F_{1}\left(a,b;c;\frac{\left(1+x\right)}{2}\right)=\sum_{j\geq 0}J_{j}\left(x|f,d\right)$
$\times\sum_{n\geq j}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}}\frac{\left(f+j\right)^{(n-j)}}{\left(n-j\right)!\left(d+f+j-1\right)^{\left(j\right)}\left(d+f+2j\right)^{\left(n-j\right)}}$
$=\sum_{j\geq 0}J_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}\times$
$\sum_{n\geq j}\frac{\left(a+j\right)^{\left(n-j\right)}\left(b+j\right)^{\left(n-j\right)}}{\left(c+j\right)^{\left(n-j\right)}\left(n-j\right)!}\frac{\left(f+j\right)^{(n-j)}}{\left(d+f+2j\right)^{\left(n-j\right)}}.$
[…]
$_{2}F_{1}\left(a,b;c;\frac{\left(1+x\right)}{2}\right)=\sum_{j\geq 0}J_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}\times$
$\sum_{s\geq 0}\frac{\left(a+j\right)^{\left(s\right)}\left(b+j\right)^{\left(s\right)}}{\left(c+j\right)^{\left(s\right)}\left(s\right)!}\frac{\left(f+j\right)^{(s)}}{\left(d+f+2j\right)^{\left(s\right)}}$
$=\sum_{j\geq 0}J_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}$
$\times_{3}F_{2}\left(a+j,b+j,f+j;c+j,d+f+2j;1\right).$
Now we apply the identity (16.4.11) of [5] and
In order to get the second identity, we proceed in a similar way, this time using polynomials $\left\{K_{n}\right\}$.
$_{2}F_{1}\left(a,b;c;(1-x\right)=\sum_{n\geq 0}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{n!\left(c\right)^{\left(n\right)}}(-1)^{n}\left(x-1\right)^{n}$
$=\sum_{n\geq 0}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}n!}(-1)^{n}\sum_{j=0}^{n}\tilde{e}_{n,j}(d,f)K_{j}(x|d,f)$
$=\sum_{j\geq 0}(-1)^{j}K_{j}(x|d,f)\sum_{n\geq j}\frac{(-1)^{n-j}\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}n!}$
$\times\frac{(-1)^{n-j}n!\left(f+j\right)^{\left(n-j\right)}}{(n-j)!\left(d+f+j-1\right)^{\left(j\right)}\left(d+f+2j\right)^{\left(n-j\right)}}$
$=\sum_{j\geq 0}(-1)^{j}K_{j}(x|d,f)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}$
$\times\sum_{n\geq j}\frac{\left(a+j\right)^{\left(n-j\right)}\left(b+j\right)^{\left(n-j\right)}\left(f+j\right)^{\left(n-j\right)}}{(n-j)!\left(c+j\right)^{\left(n-j\right)}\left(d+f+2j\right)^{\left(n-j\right)}}$
$=\sum_{j\geq 0}K_{j}(1-x|f,d)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}$
$\times~_{3}F_{2}\left(a+j,b+j,f+j;c+j,d+f+2j;1\right).$
In the last line we applied (2.18), Now it remains to replace 1x1-x by $x$ on both sides of the inequality.
2) We use (2.1) and get
$_{2}F_{1}\left(a,b;c;x\right)=(c-1)\sum_{n\geq 0}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(c-1\right)^{\left(n+1\right)}}\frac{x^{n}}{\left(1\right)^{\left(n\right)}}$
$=(c-1)\sum_{n\geq 0}\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}\frac{x^{n}}{\left(1\right)^{\left(n\right)}}\frac{1}{n!}\sum_{j=0}^{n}\binom{n}{j}\frac{(-1)^{j}}{c-1+j}$
$=\left(c-1\right)\sum_{j\geq 0}\frac{(-1)^{j}}{c-1+j}\frac{x^{j}}{j!}\sum_{n\geq j}\frac{\left(a\right)^{\left(n\right)}\left(b\right)^{\left(n\right)}}{\left(1\right)^{\left(n\right)}}\frac{x^{n-j}}{(n-j)!}$
$=\left(c-1\right)\sum_{j\geq 0}\frac{(-1)^{j}\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c-1+j\right)j!}\frac{x^{j}}{j!}\sum_{n\geq j}\frac{\left(a+j\right)^{\left(n-j\right)}\left(b+j\right)^{\left(n-j\right)}}{\left(j+1\right)^{\left(n-j\right)}}\frac{x^{n-j}}{(n-j)!}$
$=\left(c-1\right)\sum_{j\geq 0}\frac{(-1)^{j}\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}x^{j}}{\left(c-1+j\right)\left(j!\right)^{2}}~_{2}F_{1}\left(a+j,b+j;j+1;x\right)$
$=(1-x)^{1-a-b}\left(c-1\right)\sum_{j\geq 0}\frac{(-1)^{j}\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c-1+j\right)\left(j!\right)^{2}}\left(\frac{x}{1-x}\right)^{j}~_{2}F_{1}\left(1-a,1-b;j+1;x\right).$
To get the last equality, we applied the well-known Euler transformation. ∎
@@
When we use the identity (16.4.11) of [5] concerning $_{3}F_{2}\left(\begin{array}[c]{ccc}a&b&c\\ d&f&\end{array};1\right)$ we can further transform the basic expansions (3.11) and (3.15). We get then
@@
(3.31) $=\frac{\Gamma(d+f)\Gamma(c+d-a-b)}{\Gamma(d+f-a)\Gamma(d+c-b)}\sum_{j\geq 0}J_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}\times$
(3.34) $\frac{(d+f)^{\left(2j\right)}}{(d+f-a)^{\left(j\right)}(d+c-b)^{\left(j\right)}}~_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&c-b&c-f\\ c+j&d+c-b+j&\end{array};1\right)$
@@
(3.48) $=\frac{\Gamma(d+f)\Gamma(c+d-a-b)}{\Gamma(d+f-a)\Gamma(d+c-b)}\sum_{j\geq 0}K_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}\times$
(3.51) $\frac{(d+f)^{\left(2j\right)}}{(d+f-a)^{\left(j\right)}(d+c-b)^{\left(j\right)}}~_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&c-b&c-f\\ c+j&d+c-b+j&\end{array};1\right)$
@@
a) In particular setting $f\allowbreak=\allowbreak c$ and assuming that $c+d\allowbreak\neq\allowbreak 1$we get
@@
b) Also setting $d\allowbreak=f$ and assuming that $c\allowbreak=\allowbreak(a+b+1)/2$ and $2f+1>a+b$ we have
@@
$\sum_{j\geq 0}K_{j}\left(x|f,f\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}\Gamma(f+1/2+j)}{\left(\frac{(a+b+1)}{2}\right)^{\left(j\right)}\left(2f+j-1\right)^{\left(j\right)}\Gamma(\frac{(a+1+j)}{2})\Gamma(\frac{(b+1+j)}{2})\Gamma(\frac{2f+1-a+j}{2})\Gamma(\frac{2f+1-b+j}{2})}$
c) Setting $f\allowbreak=\allowbreak 2c-b-1$ and $d\allowbreak=\allowbreak 1+b+2a-2c$ and assuming that $f$ and $d$ are both positive we observe that $f+d\allowbreak=\allowbreak 2a$ and $b+f+1\allowbreak=\allowbreak 2c.$ Hence, after applying the Watson’s sum we have:
@@
$\times\frac{\Gamma(a+\frac{1}{2}+j)\Gamma(c+j)}{\Gamma(\frac{b+1+j}{2})\Gamma(c+\frac{j-1-b}{2})\Gamma(a+\frac{j+1-b}{2})\Gamma(a+1-c+\frac{j+b}{2})}$
@@
a) To get (3.56) we set $f\allowbreak=\allowbreak c$ in the either (3.11) or (3.21) then we apply Gauss summation formula and (1.1). We proceed as follows
$_{2}F_{1}\left(a,b;c;x\right)$ $=\sum_{j\geq 0}K_{j}\left(x|c,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+c+j-1\right)^{\left(j\right)}}$
$\times\frac{\Gamma\left(d+c+2j\right)\Gamma\left(d+c-a-b\right)}{\Gamma(d+c-a+j)\Gamma(d+c-b+j)}$
$=\sum_{j\geq 0}K_{j}\left(x|c,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}}{\left(c\right)^{\left(j\right)}\left(d+c+j-1\right)^{\left(j\right)}}$
$\times\frac{\left(c+d\right)^{\left(2j\right)}\Gamma(c+d)\Gamma\left(d+c-a-b\right)}{\left(d+c-a\right)^{\left(j\right)}\Gamma(d+c-a)\left(d+c-b\right)^{\left(j\right)}\Gamma\left(d+c-b\right)}.$
Now we apply the following formula true for all complex $x$ not equal to non-positive integers and $n\geq 0$ and $x\neq 1:$
$\left(x\right)^{\left(2n\right)}/\left(x+n-1\right)^{\left(n\right)}\allowbreak=\frac{\left(x\right)^{\left(n\right)}\left(x+2n-1\right)}{x-1}$
$=\frac{\Gamma(c+d)\Gamma\left(d+c-a-b\right)}{\Gamma(d+c-a)\Gamma\left(d+c-b\right)}\times$
@@
Corollary 2.

### v5

@@ 3. Hypergeometric function
We apply assertions of the Theorem 1 for $n\allowbreak=\allowbreak 2$ and $m\allowbreak=\allowbreak 1$ with $a_{1}\allowbreak=\allowbreak a,$ $a_{2}\allowbreak=\allowbreak b,$ $b_{1}\allowbreak=\allowbreak c.$ Now to transform $_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&b+j&f+j\\ c+j&d+f+2j&\end{array};1\right)$ we apply (16.4.11) of [6] and (1.1).
2) To get the last equality, we applied the well-known Euler transformation. ∎
@@
Let us notice that we can use identity (16.4.11) of [6] concerning $_{3}F_{2}\left(\begin{array}[c]{ccc}a&b&c\\ d&f&\end{array};1\right)$ in another way and get another forms of the basic expansions (3.30) and (3.37). We get then
@@
(3.67) $\times~_{3}F_{2}\left(\begin{array}[c]{ccc}f+j&d+f-b+j&d+f-a+j\\ c+d+f-a-b+j&d+f+2j&\end{array};1\right).$
@@
(3.71) $\frac{(d+f)^{\left(2j\right)}}{(d)^{\left(j\right)}(c+d+f-a-b)^{\left(j\right)}}~_{3}F_{2}\left(\begin{array}[c]{ccc}f+j&c-a&c-b\\ c+j&c+d+f-a-b+j&\end{array};1\right).$
@@
a) By setting $f\allowbreak=\allowbreak c$ and assuming that $c+d\allowbreak\neq\allowbreak 1,$ we can obtain
@@
b) Also, setting $d\allowbreak=f$ and assuming that $c\allowbreak=\allowbreak(a+b+1)/2$ and $2f+1>a+b$ we have
@@
$\sum_{j\geq 0}K_{j}\left(x|f,f\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}\Gamma(f+1/2+j)}{\left(\frac{(a+b+1)}{2}\right)^{\left(j\right)}\left(2f+j-1\right)^{\left(j\right)}\Gamma(\frac{(a+1+j)}{2})\Gamma(\frac{(b+1+j)}{2})\Gamma(\frac{2f+1-a+j}{2})\Gamma(\frac{2f+1-b+j}{2})}.$
c) Setting $f\allowbreak=\allowbreak 2c-b-1$ and $d\allowbreak=\allowbreak 1+b+2a-2c$ and assuming that $f$ and $d$ are both positive we observe that $f+d\allowbreak=\allowbreak 2a$ and $b+f+1\allowbreak=\allowbreak 2c.$ Hence, after applying Watson’s sum, we have
@@
$\times\frac{\Gamma(a+\frac{1}{2}+j)\Gamma(c+j)}{\Gamma(\frac{b+1+j}{2})\Gamma(c+\frac{j-1-b}{2})\Gamma(a+\frac{j+1-b}{2})\Gamma(a+1-c+\frac{j+b}{2})}.$
d) Let us set $a\allowbreak=\allowbreak-n,$ $d\allowbreak=\allowbreak b-c+1-n.$ We have then, after applying Pfaff–Saalschütz Balanced Sum formula
(3.73) $\left(c\right)^{\left(n\right)}~_{2}F_{1}\left(-n,b;c;x\right)=\left(c\right)^{\left(n\right)}\sum_{j=0}^{n}(-1)^{j}\binom{n}{j}\frac{\left(b\right)^{\left(j\right)}x^{j}}{\left(c\right)^{\left(j\right)}}$
$=(-1)^{n}\sum_{j=0}^{n}\frac{n!}{(n-j)!}K_{j}(x|f,d)\frac{\left(b\right)^{\left(j\right)}\left(c-b\right)^{\left(n-j\right)}\left(c-f\right)^{\left(n-j\right)}(d+f+2j-1)}{\left(d+f+j-1\right)^{\left(n+1\right)}}.$
@@
a) To get (3.72) we set $f\allowbreak=\allowbreak c$ in the either (3.30) or (3.45) then we apply Gauss summation formula and (1.1). Now we apply the following formula true for all complex $x$ not equal to non-positive integers and $n\geq 0$ and $x\neq 1:$
$\left(x\right)^{\left(2n\right)}/\left(x+n-1\right)^{\left(n\right)}\allowbreak=\frac{\left(x\right)^{\left(n\right)}\left(x+2n-1\right)}{x-1}.$
$_{2}F_{1}\left(a,b;c;x\right)==\frac{\Gamma(c+d)\Gamma\left(d+c-a-b\right)}{\Gamma(d+c-a)\Gamma\left(d+c-b\right)}\times$
@@
d) By setting $a\allowbreak=\allowbreak-n$ and $d\allowbreak=\allowbreak b-c+1-n$ we make
a balanced sum and a Pfaff–Saalschütz Balanced Sum formula ([6],16.4.3) can be applied yielding
$_{3}F_{2}\left(\begin{array}[c]{ccc}-n+j&b+j&f+j\\ c+j&\allowbreak b-c+1-n+f+2j&\end{array};1\right)$
$=\frac{\left(c-b\right)^{n-j}\left(c-f\right)^{\left(n-j\right)}}{\left(c+j\right)^{\left(n-j\right)}\left(c-b-f-j\right)^{\left(n-j\right)}}.$
Finally, we notice that $\left(c\right)^{\left(j\right)}\left(c+j\right)^{\left(n-j\right)}\allowbreak=\allowbreak\left(c\right)^{\left(n\right)}.$ Further, we notice that $\left(c-b-f-j\right)\allowbreak=\allowbreak\left(-d-f-n-j+1\right)$ and consequently that $\left(-d-n+1-f-j\right)^{\left(n-j\right)}\allowbreak$
$\allowbreak\left(-d-f-2j+1-(n-j)\right)^{\left(n-j\right)}\allowbreak=\allowbreak\left(-1\right)^{n-j}\left(d+f+2j\right)^{\left(n-j\right)},$ by (1.2). Now it is easy to notice that t and $\left(d+f+j-1\right)^{\left(j\right)}\left(d+f+2j\right)^{\left(n-j\right)}\allowbreak=\allowbreak\left(d+f+j-1\right)^{\left(n+1\right)}\allowbreak/\allowbreak\left(d+f+2j-1\right)$. ∎
Corollary 3 (Special values).
i) Let us put $x\allowbreak=\allowbreak 0$ and $x\allowbreak=\allowbreak 1$ in (3.20) and let us use 2.19. We get then the following general identities
$1=\sum_{j\geq 0}\frac{(-1)^{j}}{\left(f\right)^{\left(j\right)}}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}$
(3.76) $\times~_{n+1}F_{m+1}\left(\begin{array}[c]{cccc}a_{1}+j&\ldots&a_{n}+j&f+j\\ b_{1}+j&\ldots&b_{m}+j&\left(d+f+2j\right)\end{array};1\right).$
and
$_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};1\right)=$
$\sum_{j\geq 0}\frac{1}{\left(d\right)^{\left(j\right)}}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}$
$\times~_{n+1}F_{m+1}\left(\begin{array}[c]{cccc}a_{1}+j&\ldots&a_{n}+j&f+j\\ b_{1}+j&\ldots&b_{m}+j&\left(d+f+2j\right)\end{array};1\right).$
ii) For $n\allowbreak=\allowbreak 2$ and $m\allowbreak=\allowbreak 1$ and for $\operatorname{Re}(c-a-b)>0,$ $d>0$ the following forms are obtained by putting $x\allowbreak=\allowbreak 0$ and $x\allowbreak=\allowbreak 1$ in (3.72).
(3.77) $1$ $=\frac{\Gamma(c+d-1)\Gamma\left(c+d-a-b\right)}{\Gamma(c+d-a)\Gamma\left(c+d-b\right)}$
$\times\sum_{j\geq 0}\left(-1\right)^{j}\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}\left(c+d-1\right)^{\left(j\right)}(c+d+2j-1)}{j!\left(d+c-a\right)^{\left(j\right)}\left(d+c-b\right)^{\left(j\right)}},$
(3.78) $\frac{\Gamma(c)\Gamma(c-a-b)}{\Gamma(c-a)\Gamma(c-b)}$ $=\frac{\Gamma(c+d-1)\Gamma\left(c+d-a-b\right)}{\Gamma(c+d-a)\Gamma\left(c+d-b\right)}$
$\times\sum_{j\geq 0}\frac{\left(d\right)^{\left(j\right)}\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}\left(c+d-1\right)^{\left(j\right)}(c+d+2j-1)}{\left(j!\right)\left(c\right)^{\left(j\right)}\left(d+c-a\right)^{\left(j\right)}\left(d+c-b\right)^{\left(j\right)}}.$
iii) For and we get the following identities, true for all complex $c,b,f$ such that cbfc-b-f is not equal to a non-negative integerer less or equal to $n+1.$
(3.79) $\left(c\right)^{n}\allowbreak$ $=\allowbreak\sum_{j=0}^{n}(-1)^{j}\binom{n}{j}\frac{\left(b\right)^{\left(j\right)}\left(f\right)^{\left(j\right)}\left(c-b\right)^{\left(n-j\right)}\left(c-f\right)^{\left(n-j\right)}(c-b-f+n-2j)}{\left(c-b-f-j\right)^{\left(n+1\right)}},$
(3.80) $1$ $=\sum_{j=0}^{n}\left(-1\right)^{j}\binom{n}{j}\frac{\left(b\right)^{\left(j\right)}\left(c-f\right)^{\left(n-j\right)}\left(c-b-f+n-2j\right)}{\left(c-b-f-j\right)^{\left(n+1\right)}}.$
Proof.
Only iii) requires the proof. To get these identities, we consider (3.73) with $a\allowbreak=\allowbreak-n,$ and $d\allowbreak=\allowbreak\allowbreak b-c+1-n$ and set respectively firstly $x\allowbreak=\allowbreak 0$ and then $x\allowbreak=\allowbreak 1.$ Then recall that $\left(c\right)^{\left(n\right)}~_{2}F_{1}\left(-n,b;c;0\right)\allowbreak=\allowbreak\left(c\right)^{\left(n\right)}$ while $\left(c\right)^{\left(n\right)}~_{2}F_{1}\left(-n,b;c;1\right)\allowbreak=\allowbreak\left(c\right)^{\left(n\right)}\frac{\left(c-b\right)^{\left(n\right)}}{\left(c\right)^{\left(n\right)}}\allowbreak=\allowbreak\left(c-b\right)^{\left(n\right)}.$ Then we recall (2.19) and immediately get 3.79. To get (3.80) we have to recall that $\left(b+1-c-n\right)^{\left(j\right)}\allowbreak=\allowbreak(-1)^{j}\left(c-b+n-j\right)^{\left(j\right)}$ by (1.2). Next we notice that $\left(c-b+n-j\right)^{\left(j\right)}\left(c-b\right)^{\left(n-j\right)}\allowbreak=\allowbreak\left(c-b\right)^{\left(n\right)}$ and we cancel out this factor on both sides. ∎
Remark 6.
To make the reader realize how nontrivial the obtained identities are, let us consider identities (3.77) and (3.78) for $a\allowbreak=\allowbreak-n.$ Namely, after some elementary algebra (3.77) becomes after denoting $c+d\allowbreak=\allowbreak\alpha$ and $b\allowbreak=\allowbreak\beta,$ for all $n,$ $\alpha\neq 0,$ $\beta\neq\beta$ we get
$\frac{1}{\left(\alpha-\beta\right)^{\left(n\right)}}=\sum_{j\geq 0}\binom{n}{j}\left(1+\frac{j}{\alpha+j-1}\right)\frac{\left(\beta\right)^{\left(j\right)}\left(\alpha\right)^{\left(j\right)}}{\left(a\right)^{\left(n+j\right)}\left(a-\beta\right)^{\left(j\right)}}.$
Identity (3.78) for $a\allowbreak=\allowbreak-n$ and some algebra becomes after denoting $c-b\allowbreak=\allowbreak\alpha,$ $c\allowbreak=\allowbreak\gamma,$ $c+d-b\allowbreak=\allowbreak\beta$
$\frac{\left(\alpha\right)^{\left(n\right)}}{\left(\beta\right)^{\left(n\right)}\left(\gamma\right)^{\left(n\right)}}$ $=\sum_{j=0}^{n}(-1)^{j}\binom{n}{j}\left(1+\frac{j}{\beta+\gamma-\alpha++j-1}\right)$
$\times\frac{\left(\beta-\alpha\right)^{\left(j\right)}\left(\gamma-\alpha\right)^{\left(j\right)}(\beta+\gamma-\alpha)^{\left(j\right)}}{\left(\gamma\right)^{\left(j\right)}\left(\beta\right)^{\left(j\right)}\left(\beta+\gamma-\alpha\right)^{\left(n+j\right)}},$
which is true for $n\geq 0,$ $\gamma\neq 0,$ $\beta\neq 0,$ $\alpha\neq\beta+\gamma.$
Corollary 4 (Integrals).